Ten milliliters of phenolphthalein was transferred in a burette. What is the measured volume of the substances in the burette if its computed percentage error is 0.82%?


1
Expert's answer
2021-08-18T09:44:01-0400

Error = "\frac{10×0.82}{100}=0.082ml"


Hence measured volume = "10+0.082= 10.082ml"

Or

"10-0.082=9.918ml"